OpenStudy (anonymous):

In parallelogram DEFG, Dh=x+3, HF= 3y, GH=3x-3, and HE=5y+4. Find the values of x and y?

OpenStudy (anonymous):



OpenStudy (anonymous):

HF= HE by symmetry so 3y = 5y+4 You can solve for y Same way for x: DH= GH by symmetry

OpenStudy (anonymous):

You can set two equations since diagonal of //gram bisects each other Dh= HF =>x+3= 3y -(1) GH= HE => 3x-3=5y+4. -(2) (1)x3-(2) 3(x+3) - (3x-3)= 3(3y)-(5y+4) 12 = 4y-4 y=4 Can you solve x?